OpenStudy (anonymous):

Write the equation of a line that goes through the points (1, 4) and (2, 2)

OpenStudy (anonymous):

first get your gradient by finding rise over run (y1-y2)/(x1-x2)

OpenStudy (anonymous):

gradient is equivalent to the slope of the line

OpenStudy (anonymous):

first find the gradient then use the gradient to find the equation of line with any pair of coordinate given

OpenStudy (anonymous):

how

OpenStudy (anonymous):

A line is defined as: y = mx + b where: m = slope / gradient b = y-intercept So find m and b such that they fit the description.

OpenStudy (anonymous):

so how

OpenStudy (anonymous):

A point on a graph is denoted (x, y) such that x is a point x units along the x-axis and y is a point y units along the y-axis. If you have a point (1, 4) and (2, 2) then you know that for x = 1, y = 4, and for x = 2, y = 2. So find an equation that satisfies those conditions. Here's how you find the slope: slope = rise / run => m = dy / dx where dy = change in y-coordinates from point A to point B and dx = change in x-coordinates from point A to point B

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Do you know what Cartesian geometry is?

OpenStudy (anonymous):

so x=1 and y=4 in the first point. On a graph x coordinates will move the point left or rioght and the y coordinates will move the point up or down. The first step to making an equation is to find the slope (How much the point has gone up or down divided by how much the point has gone left or right.)

OpenStudy (anonymous):

So from X=1 to X=2 we move right +1 on the graph. The run is +1

OpenStudy (anonymous):

from Y=4 to Y=2 we move down -2 so the rise is -2.

OpenStudy (anonymous):

Put them together for the slope. Rise/Run = 1/-2 or -1/2

OpenStudy (anonymous):

Now that we know the slope we need to find the y-intercept. Where the line crosses the y axis

OpenStudy (anonymous):

I did the slope backwards :3 It's actually -2/1 or just -2

OpenStudy (anonymous):

Y=2 at X=1 so what will Y be at X=0? Use the slope and trace the line back from (1,4). You should reach (0,6).

OpenStudy (anonymous):

Now to make the equation. y = mx + b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

m=-2 and b=+6 so your equation comes out to be y = -2x + 6

OpenStudy (anonymous):

And that's how math gud.

OpenStudy (anonymous):